### v2

# ParisLuco3D: A high-quality target dataset for domain generalization of LiDAR perception

Abstract
LiDAR is an essential sensor for autonomous driving by collecting precise geometric information regarding a scene. As the performance of various LiDAR perception tasks has improved, generalizations to new environments and sensors has emerged to test these optimized models in real-world conditions. Unfortunately, the various annotation strategies of data providers complicate the computation of cross-domain performances.
This paper provides a novel dataset, ParisLuco3D, specifically designed for cross-domain evaluation to make it easier to evaluate the performance utilizing various source datasets. Alongside the dataset, online benchmarks for LiDAR semantic segmentation, LiDAR object detection, and LiDAR tracking are provided to ensure a fair comparison across methods.
The ParisLuco3D dataset, evaluation scripts, and links to benchmarks can be found at the following website:
https://npm3d.fr/parisluco3d
I Introduction
LiDAR-based perception for autonomous driving applications has become increasingly popular in the last few years. LiDAR provides reliable and precise geometric information and is a useful addition to typical camera-based systems. The various LiDAR perception tasks have gained access to a growing number of open-source datasets. This large amount of data, and the thorough efforts of the community, have led to very good performance, notably by Cylinder3D [1] for LiDAR Semantic Segmentation (LSS) and CenterPoint [2] for LiDAR Object Detection (LOD).
Consequently, new tasks for evaluating the robustness of these methods in new scenarios and environments have emerged. Here, we focus on methods that deal with domain generalization, which have become an increasing focus in LiDAR perception. These methods involve confronting a model trained on a specific domain, with a new domain at the time of inference. In practice, a subset of datasets are used for training, and other datasets expected to be acquired elsewhere are used for testing.
While open-source datasets display a large variety of scenes and sensor setups, crucial discrepancies have been identified among label sets [3, 4]. Due to this label difference, it is not possible to leverage current online benchmarks to ensure fair comparisons across methods.
Figure 1: A LiDAR scan of our ParisLuco3D dataset with ground truth annotation for semantic segmentation and object detection.
Name Year # scans train for LSS # scans validation for LSS LSS LOD Tracking Location Scene type
Sensor resolution (${}^{\circ}$)
# beams
vertical
horizont.
KITTI [5] 2012 N/A N/A $\times$ $\sim$ $\times$
Germany
suburban
0.4 0.08 64
SemanticKITTI [6] 2019 19k 4071 $\times$
Germany
suburban 0.4 0.08 64
nuScenes [7] 2020 29k 6019
US+Singapore
urban 1.33 0.32 32
SemanticPOSS [8] 2020 N/A ${}^{\dagger}$$\times$ $\times$
China
campus 0.33-6 0.2 40
Waymo [9] 2020 24k 5976
US
urban 0.31 0.16 64
ONCE [10] 2021 N/A N/A $\times$ $\checkmark$ $\times$
China
sub.+urban
0.33-6 0.2 40
Pandaset [11] 2021 N/A ${}^{\dagger}$
US
suburban 0.17-5 0.2 64
HelixNet [12] 2022 49k 8179 $\times$ $\times$
France
sub.+urban
0.4 0.08 64
KITTI-360 [13] 2023 60k 15204 $\sim$ $\times$
Germany
suburban 0.4 0.08 64
ParisLuco3D (Ours) 2023 N/A ${}^{\dagger}$ France urban 1.33 0.16 32
TABLE I: Summary of the various existing LiDAR perception datasets. ${}^{\dagger}$size of the full dataset, as there is no designated split. KITTI provides LOD annotation only for a subset of the point cloud. KITTI-360 provides the LSS annotation on the accumulated and subsampled point cloud rather than on each scan (LSS = LiDAR Semantic Segmentation and LOD = LiDAR Object Detection).
In this paper, we propose a new dataset that is specifically designed for domain generalization evaluation, with annotations that have been mapped to existing standard datasets to avoid the need for comprise in quality evaluations of the domain generalization methods. Since this dataset is focused only on cross-domain generalization evaluation, its annotations are not made available to the public.
We make the following contributions:
  • The release of a novel dataset, in addition to online benchmarks, to compute cross-domain performances fairly.
  • A thorough overview of the generalization capabilities of current state-of-the art architecture for both semantic segmentation and object detection in order to provide the community with a baseline for these tasks.
II Related Work
II-A LiDAR Datasets in autonomous driving
The tasks that we focus on for LiDAR perception are LiDAR Semantic Segmentation (LSS), LiDAR Object Detection (LOD) and Tracking.
A summary of the main datasets along with the tasks they are annotated for, sensor information, and the amount of scans available is presented in Table I.
The two most common perception tasks, measured by the number of annotations available or the number of published works in these fields, are semantic segmentation and object detection. Accordingly, these tasks are the focus of the remainder of this work. Recently, some large scale datasets have been released, such as HelixNet [12], DurLAR [14] or Argoverse 2 [15] but due to their recency or lack of annotations, they are not yet widely used in the LSS and LOD literature.
II-A1 Datasets for LiDAR semantic segmentation
The first dataset released for LSS was SemanticKITTI [6] in 2019, which was derived from the KITTI [5] dataset. It was acquired in a German suburb. Due to its size and seniority, it is the reference dataset for LiDAR semantic segmentation. Its annotations are evenly split between road users and background classes.
Thereafter, SemanticKITTI was expanded into KITTI-360 [13], which includes data on the same suburb but on a larger scale.
Subsequently, a few object detection datasets were expanded to incorporate semantic segmentation annotation, such as nuScenes [7] and Waymo [9]. They were acquired in cities and suburbs in the US and Asia. Contrary to SemanticKITTI, nuScenes focuses its annotations on vehicles with a limited number of background classes.
Finally, a few additional targeted datasets were released, such as SemanticPOSS [8], which proposed a new setup by providing a dataset acquired in a student campus, and Pandaset [11], which has a double-sensor setup.
II-A2 Datasets for LiDAR object detection
In 2014, KITTI [5] was released and ushered in a new range of experimental data for LOD. While there are several downsides in its annotation, such as the low number of classes and the limited annotated field of view, it has remained an important benchmark for detection due to the quality of its annotations and its high-density point clouds.
The nuScenes [7] dataset was subsequently released with full 360-degrees annotations and an exhaustive list of annotated classes. Even though the number of available scans in this dataset is higher than those in KITTI, they are much sparser due to the frequency of the sensor and its resolution.
The Waymo [9] dataset is the one with the most 3D annotations, along with a high number of high-density point clouds. Nevertheless, it only includes coarse class information, whereas other datasets have a finer label space.
Finally, new datasets were released in an attempt to diversify the environments and LiDAR sensors used for detection, such as the ONCE [10] dataset. This dataset uses a 40-beam LiDAR sensor, therby resulting in an atypical point cloud density distribution.
II-B Domain generalization for LiDAR perception
Domain generalization was long reserved for 2D perception [16], as domain adaptation was the main focus of 3D research. The difference between these fields is that domain generalization does not utilize any information before processing the new domain. But models have been found to be very sensitive to domain gaps in datasets such as differences in LiDAR density or object sizes [17].
Recently, a few studies on domain generalization of LiDAR perception have emerged, particularly for object detection [18] by adversarial augmentation and for semantic segmentation [3, 19, 20] using domain alignment by identifying a canonical representation either in the feature space or in the Euclidean space. Additionally, MDT3D [21] is focused on multi-source domain generalization.
All the recent studies highlight the difficulties of performing cross-domain evaluation, as annotations can and do differ from one dataset to another. Specifically, they have to resort to studying either the intersection of the label set [20], usually on very few labels, or a remapping to a common and coarser label set [3].
ParisLuco3D provides a new annotation set that has been mapped to standard LiDAR perception datasets. Cross-domain evaluation can be performed fairly without any compromise regarding the fineness of the label set. Furthermore, while this dataset could be considered small, as it is only used for evaluation, it is on par or even bigger than most validation sets (Table I).
III Dataset presentation
III-A Acquisition and data generation
III-A1 Acquisition
Figure 2: Trajectory of our dataset (2.1 km around the Luxembourg Garden in Paris) overlaid on a Google Satellite Image.
The acquisition of the ParisLuco3D dataset was performed in Paris in November 2015 over a distance of 2.1 km. The trajectory, depicted in Figure 2, is around the Luxembourg Garden in the heart of the city of Paris. Figure 3 depicts point clouds by accumulating the scans and illustrates the diversity of scenes and contexts in our dataset.
The LiDAR sensor used was a Velodyne HDL32 positioned vertically on a pole that was attached to the roof of a Citröen Jumper vehicle. Due to its placement and size of the vehicle, the sensor was approximately 3.70 m from the ground (a height that is different from what is usually available in other LiDAR datasets in autonomous driving).
III-A2 Data generation
The ParisLuco3D dataset consists of 7501 scans in the form of 3D point clouds (a scan is a 360-degree horizontal rotation of the LiDAR sensor rotating at 10Hz). The data are available in raw format with the following information: x, y, z, timestamp, intensity, and laser_index. Timestamp is the time per point provided by the LiDAR sensor, synchronized with a GPS sensor. Intensity is the strength of the return signal, which represents 256 calibrated reflectivity values (diffuse reflectors are represented with values ranging from 00 to $100$ and retro-reflectors from 101–255) [^1] https://tinyurl.com/ycxm2d7a. Laser_index refers to the number of the laser that was fired (32 lasers in the Velodyne HDL32). Timestamp and laser_index, which are often not available in other datasets [6, 7, 9, 8, 11], can enable the testing of specific methods, such as the Helix4D [12], which exploits the timestamp.
Figure 3: Illustrations depicting the diversity of scenes and quality of annotation in the ParisLuco3D dataset (left: point clouds by accumulating the scans of our dataset colorized with labels33footnotemark: 3; right: images taken from Google Street View).
This raw data do not take into account the movement of the vehicle during a scan. To estimate this movement, we used CT-ICP [22] LiDAR odometry, which enables the calculation of the deformation of the scans. Therefore, we also provide motion-corrected scans with the same information as raw scans: x, y, z, timestamp, intensity, and laser_index. All data annotations (semantic class per points, bounding boxes of objects, and tracking of objects) were made on these motion-corrected scans.
Finally, to enable the use of methods that utilize data sequentiality (i.e., past scans), we again used CT-ICP to compute a pose for each scan (motion-corrected) in a world frame corresponding to the frame of the first scan of the sequence. The poses are provided in KITTI format [5] and correspond precisely to the pose of the vehicle in the middle of a scan acquisition (i.e., when the LiDAR lasers fire forward like the KITTI convention).
The sequence is composed of a loop, with an overlap of approximately 150 m between the start and end of the acquisition. A registration by ICP on the overlap zone and a graph SLAM solver [23] enabled us to calculate the final poses with loop closure (Figure 2).
III-B Label shift and annotation process
III-B1 Label shift
A major issue when conducting cross-dataset evaluation is the difference in label set definitions. There are three sources of differences, which we illustrate for the semantic segmentation case.
First, certain objects are annotated in one dataset but not in another. An example of this is the traffic-cone annotation found in nuScenes but not in SemanticKITTI.
Second, there are granularity differences. For example, nuScenes defines manmade, whereas SemanticKITTI distinguishes building, pole, and sign. Similarly, Waymo uses a single Vehicle class instead of making a distinction between cars, trucks, and buses.
Finally, there is a phenomenon we call the label shift, which represents different objects under similar labels. An example of label shift is the definitions of road and sidewalk in nuScenes and SemanticKITTI. In one dataset, bike lanes are part of the road label, whereas in the other dataset they part of the sidewalk label.
Furthermore, label shift appears and results in change for bounding box dimensions. For example, the Waymo dataset does not include large objects carried by pedestrians in its bounding boxes and does not annotate any parked bicycles.
III-B2 Annotation process
In order to guarantee the quality of the annotations in this dataset, we relied on expert annotators working in the field of 3D point cloud processing who have experience in working with existing datasets. The dataset was triple-checked in order to identify any errors.
Semantic annotation was done on the dense point clouds accumulated from motion-corrected scans. The annotation of objects for detection and tracking was done scan by scan by exploiting neighboring scans when the objects were occluded with a bounding box interpolated over time. The software used for the semantic segmentation annotation was point labeler [6], while the one for object detection was labelCloud [24], which are both open-source.
III-C Labels for semantic segmentation
[…]
SemanticKITTI [6] 19 N/A 10 19
nuScenes [7] 16 10 N/A 16
Pandaset [11] 37 8 8 13
Waymo [9] 22 15 11 20
KITTI-360 [13] 18 15 13 16
SemanticPOSS [8] 13 11 6 13
HelixNet [12] 9 7 6 7
ParisLuco3D 45 19 16 N/A
TABLE II: Number of labels of each LSS dataset and the size of the intersection of their label set with SemanticKITTI (SK) and nuScenes (NS).
In Table II, we summarize the number of labels from the various datasets and indicate how many of these labels are intersecting with the label sets of nuScenes and SemanticKITTI. We chose these two datasets because they are the commonly used training sets. When cross-domain evaluation is performed, we see a significant reduction in the number of labels for evaluation and a loss of fineness for prediction.
For ParisLuco3D, we wanted to define a label set that could be easily mapped to nuScenes and SemanticKITTI, as they are the two reference training sets. As such, the annotation details of nuScenes and SemanticKITTI were dissected to understand which objects are encapsulated in each label to avoid a label shift between ParisLuco3D and these datasets. This results in fine annotation, thereby avoiding any ambiguity in the remapping.
There are then 45 labels for semantic segmentation: car, bicycle, bicyclist, bus, motorcycle, motorcyclist, scooter, truck, construction-vehicle, trailer, person, road, bus-lane, bike-lane, parking, road-marking, zebra-crosswalk, roundabout, sidewalk, central-median, building, fence, pole, traffic-sign, bus-stop, traffic-light, light-pole, bike-rack, parking-entrance, metro-entrance, vegetation, trunk, vegetation-fence, terrain, temporary-barrier, pedestrian-post, garbage-can, garbage-container, bike-post, bench, ad-spot, restaurant-terrace, road-post, traffic-cone, and other-object.
These labels cover the usual autonomous driving labels as well as the more specific features of specific cases of the Paris landscape, such as metro entrances or bar terraces.
Due to the nature of the scene, we observe a large quantity of pedestrians and buses. The details of the label distribution are presented in Figure 4.
Figure 4: Distribution of labels in ParisLuco3D.
III-D Labels for object detection
As is the case for most autonomous driving datasets, we mainly focused on road agents and cover them all with our 11 labels. These are car, bus, truck, trailer, bicycle, bicyclist, motorcycle, motorcyclist, scooter, scootercyclist, and pedestrian.
We took into account the specificities of the common classes between datasets when choosing our labels and the level of granularity to ease the mapping when working in a cross-dataset setting, which we believe to be relevant for the generalization task. We also seek to be as fine-grained as possible for road users, and base our fine-grained LOD classes on the nuScenes dataset, which are easily mappable to the other datasets.
Whenever possible, we also distinguish between static objects and dynamic ones, such as bicycle and bicyclist. Further, due to the high redundancy between LiDAR scans, we only annotated one out of ten scans with bounding box annotations, while other scans were also used to help guide the annotation process. As depicted in Table III, we have the highest number of pedestrians per annotated LiDAR scan compared to other datasets, which are notoriously difficult to detect at long ranges.
Name Classes 3D Boxes
Instances per scan
car bicycle pedestrian
KITTI [5] 3 18k 3.8 0.2 0.6
nuScenes[7] $*{}^{*}$ 149k 11 0.3 5.7
Waymo [9] 3 1.8M 29.8 0.3 13.5
ONCE [10] 5 90k 19.8 6.3 2.9
ParisLuco3D ${}^{**}$ 64k 14.7 6.4 56.3
TABLE III: Comparison between 3D object detection datasets. All values are computed for the validation splits of the datasets. $*{}^{*}$he 23 original classes of nuScenes are grouped during evaluation into 10 due to a strong similarity in a few fine classes—for example, standing pedestrians and sitting pedestrians. ${}^{**}$hile we annotate 11 different classes, we evaluate 7 of these classes for models trained on nuScenes and 5 for those trained on ONCE, using the intersection of classes between these datasets and ours.
Due to the natural occlusion phenomenon that occurs during traversal of the environment, certain annotated objects may not contain any points. However, if during the annotation process we became aware of the presence of an object, for example by aggregating sequential point clouds, we maintained its annotation in order to facilitate both annotation and the tracking task.
While we annotated bounding boxes to their maximum range, certain filters are applied while evaluating them in order to ensure a fair evaluation process. First, only boxes whose center was within 50 m of the sensor were retained. Second, since many boxes were kept even if they did not contain any points, we did not expect any model to predict them. In practice, we filter out boxes that contain five points or less for object detection evaluation.
For a given 3D scan containing $N$ annotated objects, bounding box annotations are represented in the following manner:
$\{b_{i}=(cx_{i},cy_{i},cz_{i},w_{i},l_{i},h_{i},\theta_{i})\}_{i\in[1,N]}$
where $cx_{i}$, $cy_{i}$, and $cz_{i}$ represent the center of the bounding box annotation along the $x$, $y$, and $z$ coordinates; $w_{i}$, $l_{i}$, and $h_{i}$ represent the width, length, and height; and $\theta_{i}$ represents the heading of the bounding box.
III-E Labels for tracking
Every element within the object detection annotations was tracked. This implies that there are also 11 classes for tracking, and one of out 10 scans were annotated.
While detection computes scores for boxes populated with more than five points, objects that drop below five points due to occlusion are still tracked.
IV Overview of LiDAR Domain Generalization on ParisLuco3D
To illustrate the importance of providing a new dataset that specifically targets domain generalization, we benchmarked current state-of-the-art methods of semantic segmentation and object detection, we used ParisLuco3D as the target dataset to avoid the need to compromise on the label set at the time of evaluation.
IV-A LiDAR semantic segmentation
IV-A1 Related work
LSS builds upon the typical 2D convolutionnal neural networks. Earlier approaches projected the 3D point cloud into a 2D representation, such as a range projection [25] or a bird’s-eye-view (BEV) projection [26]. While these methods were extremely fast, their performance was not satisfactory
In parallel, point-based architectures were developed by taking the time to redefine the convolution [27].
The highest performing methods are the sparse voxel-based [28] models, which restructure the point cloud into a regular 3D grid and apply 3D convolution to it [1].
Recently, a few methods have begun to exploit the acquisition pattern of typical LiDAR sensors to achieve a high inference speed with a limited decrease in performance [12].
IV-A2 Experiments
To compute domain generalization performance, two datasets are required. One is used as the training set, called the source set, while the other is used as the test set, called the target set. The target set is supposed to have never been seen at any point in the training process. As source sets, we utilized two datasets considered to be standard: SemanticKITTI and nuScenes. They are considered standard due to their size and seniority. These two source datasets evaluate different domain gaps to the target dataset ParisLuco3D:
  • SemanticKITTI: LiDAR sensor (VelodyneHDL64 $\rightarrow$ VelodyneHDL32 with different positioning in height), environment (suburban $\rightarrow$ urban), country (Germany $\rightarrow$ France)
  • nuScenes: LiDAR sensor (exactly the same LiDAR sensor but different positioning in height), environment (urban $\rightarrow$ urban), country (US+Singapore $\rightarrow$ France).
We computed the generalization results for seven different neural architectures from SemanticKITTI to ParisLuco3D and from nuScenes to ParisLuco3D (Table IV). These different models were selected due to their representativeness of the various input types (projection-based, point-based and voxel-based), the availability of the source code, and the quality of their source-to-source results. All methods were trained with the respective original label sets of SemanticKITTI and nuScenes, and evaluated on that basis.
To date, very few domain generalization methods for LiDAR point clouds have been released. The most notable one [3] has not made its code available and, thus, we did not benchmark it. Nonetheless, we followed their protocol to create a generalization baseline by using IBN-Net [29] and RayDrop [30] in addition to SRUNet. IBN-Net relies on modifying the typical convolution block by incorporating instance normalization to reduce the statistical overfitting to the dataset. RayDrop randomly discards the acquisition ring to simulate lower resolution scans in order to increase the variance of the point distribution within a scan.
The metric used for the computation of the semantic segmentation performance is the Intersection over Union (IoU) for each class and its mean (mIoU). It is measured at a distance of up to 50 meters away from the sensor, which is similar to SemanticKITTI evaluation.
$SK\rightarrow PL$ $NS\rightarrow PL$*
[…]
CENet [25] Range 58.8 21.6 69.1 31.6
PolarSeg [26] BEV 61.8 9.6 71.4 11.3
KPConv [27] Point 61.8 20.3 64.2 22.9
SRUNet [28] Voxel 63.2 30.7 69.3 37.4
SPVCNN [31]
Point+Voxel
63.4 28.9 66.8 38.7
Helix4D [12] 4D Point 63.9 18.3 69.3 19.2
Cylinder3D [1]
Cyl. voxel
64.9 23.0 74.8 25.5
SRUNet [28]
+RayDrop [30]
Voxel 61.7 35.8 66.4 31.6
SRUNet [28]
+IBN-Net [29]
Voxel 64.9 28.0 67.3 39.7
TABLE IV: Generalization baseline when trained on SemanticKITTI (SK) or nuScenes (NS), and evaluated on ParisLuco3D (PL). *When nuScenes is the source, the intensity channel of the LiDAR sensor is used.
IV-A3 Results and analysis
Table IV first shows that good performance of a neural architecture in source-to-source does not imply good performance in target at all. In addition, generalization methods (Raydrop, IBN-Net) do not systematically allow a performance gain on the target. Table IV indicates that the best performing generalization neural architecture is the voxel-based one—namely SRUNet [28]. This is why we selected this architecture to test the IBN-Net and RayDrop generalization methods. But overall, the performance of all architectures on ParisLuco3D is disappointing compared to their performance on the training set.
For SemanticKITTI to ParisLuco3D, all architectures were retrained without the intensity channel. As the sensors used in the two datasets varied, the intensity decreased the generalization performance. In Table V, we illustrate this decrease in performance in several cases.
[…]
$SK\rightarrow PL$ $NS\rightarrow PL$
[…]
SRUNet [28]
with intensity
66.6 27.9 69.3 37.4
SRUNet [28]
without intensity
63.2 30.7 66.3 32.3
Cylinder3D [1]
with intensity
70.4 2.7 74.8 25.5
Cylinder3D [1]
without intensity
64.9 23.0 70.2 17.1
TABLE V: Impact of the intensity channel on generalization performance for semantic segmentation task.
With regards to domain generalization methods, we can observe that IBN-Net is useful for nuScenes towards ParisLuco3D and detrimental for SemanticKITTI. It stems for the increased similarity in instance statistics between nuScenes and ParisLuco3D due to the intensity channel. Conversely, RayDrop is useful in the SemanticKITTI towards ParisLuco3D case. Using RayDrop helps the model learn to segment a larger variety of lower sensor resolution. Due to the inability of methods to faithfully increase the resolution, it is detrimental when the original resolution is lower than the target resolution, which is the case for nuScenes (LiDAR rotating at 20Hz) towards ParisLuco3D (same LiDAR rotating at 10Hz).
IV-B LiDAR object detection
IV-B1 Related work
LOD predicts bounding boxes by applying local shape feature extractors and using convolutional filters similarly to 2D object detection. Works such as SECOND [32] and PointPillars [33] sought to increase the speed of using these convolutions to enable real-time applications.
However, point-based methods also exist. For example, PointRCNN [34] obtains point-wise multi-range features to optimize predictions.
The best performing methods typically use two-steps processes of initial proposals and refinement using features interpolated from keypoints [35, 2]. Recently, methods have been exploring predictions based on object centers [2] as well as transformer-based architectures [36].
IV-B2 Experiments
We established a benchmark for testing models on our ParisLuco3D dataset. As we aimed to test the domain generalization performance of detection models, we trained them on two standard source datasets: ONCE and nuScenes. Then, we evaluated them on ParisLuco3D, using the latter as a test set. Therefore, ParisLuco3D scans and annotations were not utilized during the training process. We chose ONCE due to its size and its differences in LiDAR sensors, a 40-beam sensor. The source dataset ONCE evaluates different domain gaps to the target dataset ParisLuco3D:
  • ONCE: LiDAR sensor (40-beam LiDAR $\rightarrow$ VelodyneHDL32 with different positioning in height), environment (sub.+urban $\rightarrow$ urban), country (China $\rightarrow$ France)
We evaluated the generalization ability of five different 3D object detection models across the various choices of source datasets: SECOND [32], PointRCNN [34], PointPillars [33], PV-RCNN [35], and CenterPoint [2]. These models were selected due to their availability and their performance across different 3D object detection benchmarks.
We also implement the same methods for domain generalization as previously used for semantic segmentation—that is instance normalization IBN-Net [29] and ray-dropping data augmentations inspired by Theodose et al. [30].
The predictions were evaluated by considering those with low overlap with ground-truth bounding boxes as false positives. We used IoU thresholds specific to each class for this—specifically 0.7 for four-wheel vehicles, 0.5 for two-wheel vehicles, and 0.3 for pedestrians. Furthermore, we use the average precision (AP) metric, which is commonly used for the 3D object detection task, and its mean across all classes (mAP). The average precision for a given class is defined as the precision integrated over a fixed range of recall values. We selected 50 recall values ranging between 0.02 and 1, following the protocol of ONCE.
IV-B3 Results & analysis
$ON\rightarrow PL$ $NS\rightarrow PL$
[…]
PointRCNN [34] 28.6 8.3 18.4 10.6
PointPillars [33] 45.5 12.4 35.3 10.9
PV-RCNN [35] 52.4 10.6 29.6 17.3
SECOND [32] 54.0 9.9 38.6 12.9
CenterPoint [2] 59.5 8.5 35.0 13.4
PointPillars [33]
+RayDrop [30]
39.4 13.3 29.8 9.5
PointPillars [33]
+IBN-Net [29]
45.0 11.1 35.9 10.3
TABLE VI: Generalization baseline for LOD when trained on ONCE (ON) and nuScenes (NS) and evaluated on ParisLuco3D (PL). The intensity channel is not used.
We present our benchmark results in Table VI in terms of mAP. Our models were trained without the intensity channel, as we emphasize in Table VII how intensity hinders the generalization ability of models. In Table VI, despite its relatively low source mAP, the PointPillars models tends to generalize better on ParisLuco3D when trained on ONCE. We believe that this is because it contains a lower number of parameters, and so it tends to be less susceptible to overfitting. On the other hand, PV-RCNN works best when training using nuScenes, as the model is natively capable of capturing more fine-grained details, which is necessary when working on the low resolution of nuScenes.
Table VI highlights the fact both RayDrop and IBN-Net are insufficient generalization baselines and their impact may depend on the source dataset used. For example, using the RayDrop augmentation for a PointPillars model trained on ONCE when the source LiDAR density is higher than that of ParisLuco3D tends to help the generalization ability of this model ($+0.9$ mAP). However, the opposite occurs when we use RayDrop with nuScenes ($-1.2$ mAP). RayDrop increase the generalization for most classes for our PointPillars model trained on ONCE, making it the model with the highest mAP on ParisLuco3D despite having one of the lowest mAP performances on its own source datasets. We believe this highlights the fact that aligning the LiDAR resolutions while taking advantage of large datasets can be a strong generalization method. However, as shown using nuScenes as a training set, it is not a one size fits all method and is dependant on the difference in LiDAR sensor resolutions between a source and a target dataset, which is normally unknown in the conditions of the domain generalization task.
$ON\rightarrow PL$ $NS\rightarrow PL$
[…]
SECOND [32]
with intensity
54.1 7.1 45.9 8.9
SECOND [32]
without intensity
54.0 9.9 42.1 8.0
CenterPoint [2]
with intensity
62.5 6.5 41.1 10.5
CenterPoint [2]
without intensity
59.5 8.5 38.7 15.6
PointPillars [33]
with intensity
45.0 9.2 40.9 2.5
PointPillars [33]
without intensity
45.5 12.4 37.7 12.7
TABLE VII: Impact of the intensity channel on generalization performance for the detection task.
In Table VII, we trained SECOND, CenterPoint, and PointPillars on the ONCE and nuScenes datasets both with and without intensity. It is evident that unlike in semantic segmentation, using intensity for object detection when training on nuScenes actually hinders the final performance by a large amount, ranging from a +0.9 to -10.2 mAP difference on ParisLuco3D compared to models trained without intensity (yet nuScenes and ParisLuco3D have the same LiDAR sensor). We believe this is due to a difference in the intensity distributions of objects, which has been shown to be a major factor of model overfitting for LOD [37].
Overall, we found that both LSS and LOD models suffer from large accuracy drops when transferred to our ParisLuco3D dataset, thereby highlighting the need for novel neural architectures and specific 3D generalization methods. The intensity channel of LiDAR is an input which is a source of overfitting in source-to-source and requires further research to be better exploited in domain generalization (even with the same LiDAR sensor as target).
V Online benchmark
To ensure a fair measurement of the domain generalization performance on this dataset, only the raw dataset is released and not the ground-truth labels. For qualitative comparisons of methods, five scans have been released alongside their annotations.
We then establish online benchmarks to test the domain generalization for three LiDAR perception tasks: semantic segmentation, object detection, and object tracking.
All details for online benchmarks are available on the project website: https://npm3d.fr/parisluco3d
We encourage authors to avoid typical benchmark optimization techniques in order to fairly judge generalization performance. Following [38], the number of submissions is limited.
VI Conclusion
In this paper, we proposed a new dataset for LiDAR perception. This dataset stands out from other already released datasets due to the fineness of the annotations, and its goal of domain generalization evaluation.
We also benchmarked different neural architectures and different generalization methods for LiDAR semantic segmentation and LiDAR object detection, thereby demonstrating that the most recent and efficient architectures are not the most robust and that generalization methods have mixed results for now.
We hope that this dataset will enable visibility in our community regarding the emergence of new 3D perception methods that generalize well, thereby allowing 3D perception tasks to be used in real-world conditions.
Acknowledgments
We would like to thank Hassan Bouchiba, who mainly contributed to the acquisition of this dataset back in 2015.